OpenStudy (anonymous):

For a 2x2 matrix, A = [a_ij], whose elements are given by a_ij = i/j Write the value of a_12 I put the i = 1 and j = 2, then what? I'm forgetting things as I have an exam tomorrow :(

ganeshie8 (ganeshie8):

yes and have a look at the given rule : a_ij = i/j

ganeshie8 (ganeshie8):

first index represents the ROW number second index represents the COLUMN number

ganeshie8 (ganeshie8):

a_ij is the element at i'th row and j'th column

OpenStudy (kainui):

In my mind I like to pair it with the idea of how in algebra we sometimes call slope "rise over run" just as a useful trick to help me remember what order they go in. Similarly here, you're "falling" for the first index and "running" for the second index.

OpenStudy (anonymous):

So the answer is simply 1/2?

ganeshie8 (ganeshie8):

\[\large \left[ \begin{array}{ccc} 1/1 & \color{red}{1/2} & 1/3 \\ 2/1 & 2/2 & 2/3 \\ 3/1 & 3/2 & 3/3 \end{array} \right]\]

OpenStudy (kainui):

In fact, the matrix is completely meaningless to this question. a_ij = i/j means a_12=1/2

ganeshie8 (ganeshie8):

thats the element at first row, and second column ^^

OpenStudy (anonymous):

Oh Okay that's it? that's fairly simple :D

ganeshie8 (ganeshie8):

thats it ! we're just doing the new postman job :P

OpenStudy (anonymous):

Wish I could give you both a medal @Kainui @ganeshie8 You guys are gonna help me pass my exam :')

ganeshie8 (ganeshie8):

locating the houses using the addresses